# A Compressive-Expressive Communication Framework for Compositional Representations

### v3

@@ Appendix D Recovering compositional structure from partially observed data
Then Prop. 3.1 states the following:
@@
1. For $i<C$ and any choice of $M,\epsilon>0$, the restrictions on $i$-token submessages of $S(x)$ by requiring $S\in\mathcal{S}_{full}^{M}$ are more restrictive than those imposed by requiring $S\in\mathcal{S}_{final}^{\epsilon}$, i.e. we have $\mathcal{R}_{[i]}(\mathcal{S}_{full}^{M})\subset\mathcal{R}_{[i]}(\mathcal{S}_{final}^{\epsilon})$.
@@
Item 1 is direct, as no restriction on intermediate submessages $S(x)_{[i]}=\mathcal{R}_{[i]}(S(x))$ by the requirement $S\in\mathcal{S}_{final}^{\epsilon}$.
@@ Appendix F Regularization to counteract Holistic Encodings
This corresponds to the balance between compositional and parsimonious communication, a general theme in classical IL literature (see e.g. [27, 29]). It seems useful to specify more precise definitions in our generative setting, in order to help justify our architecture choices.
@@
    In §F.1 we set rigorous definitions of Holistic Encodings within the setting of §2.1 and §2.2.
@@ F.1 Definitions of Holistic and Compositional encodings, and role of the compressibility condition
In the classical IL setup [27, 29], a holistic system of communication is defined as an encoding $\mathcal{G}\to\mathcal{M}$ that maps $\mathcal{G}\ni\mathbf{G}\mapsto\mathbf{G^{\prime}}\in\mathcal{M}$ injectively but disregards the compositional structure of the "language" to be encoded, i.e. in our case, without respecting the structure of $\mathbf{G}$. In our generation task formalization, a difference is that the relevant encoding map now factors through the generation map as $\mathcal{G}\stackrel{{\scriptstyle\mathsf{GenX}}}{{\longrightarrow}}\mathcal{X}\stackrel{{\scriptstyle S}}{{\longrightarrow}}\mathcal{M}$. Now the generation map $\mathsf{GenX}$ is considered fixed, thus the definition refers to only $S$ and is the following:
@@
Note that, as in [27, 29], Definition F.1 of holistic encodings refers to "respecting compositional structure", which requires another definition. The definition of compositional encodings crucially depends on how complex we need/want to assume $\mathbf{G}$ to be. In this paper, for the sake of concreteness, we restricted to the simple case of $\mathbf{G}$ being formed as an element of a product space of finite cardinality, i.e., an $n$-ple of independent factors $\mathbf{G}=[G_{1},\dots,G_{n}]$ in which each $G_{i}$ can take finitely many values. We thus use a straightforward toy definition of compositional encodings, requiring that each of the $G_{i}$ is separately encoded as a string and these strings are concatenated together to give the message that encodes $\mathbf{G}$:
@@ F.3 Encodings that optimize efficiency under expressivity over $\mathcal{D}_{train}$ are holistic
[…]
@@
    then ()($\star$)( ⋆ ) becomes sharp, again guaranteeing $C_{0}>C_{\star}$ as desired. This could be probable to happen in our setting, and becomes more likely for large $|V|>1/p$ for which (25) fails, given that the $N_{i}$ are supposed to be unknown, so it is likely that for large $|V|>1/p$ for a few of the $N_{i}$ the gaps in (26) are nontrivial. For example, (26) holds true for $|V|>1/p\geq 10$ if, say, at least two of the $G_{i}$ are binary factors, so that $N_{i}=2$.
@@ F.4 Possible choices of encoded message regularization
    Background and standard approach. Introduced by [57] for use in policy-based reinforcement learning, entropy maximization is known to enforce exploration in deterministic policies and avoid early convergence to single output choices for learned policies. Specifically, in a learning setup with input $x$, output $y$ and policy parameters $W$, the authors define the following estimator for fixed output value $\xi$:
@@
    $-\ln Pr(y=\xi|W,x)=-\sum_{i\in U}g_{i}(\xi_{i},w^{i},x^{i})=h(y,W,X).$ (31)
    In fact, this method can be extended onto sequence prediction models, and even IL, such as in [42].
@@
    $(\exists v\in V),\quad Pr(m_{t}\mid x,W,m_{[t-1]})\approx\begin{dcases}1,&\text{ if }m_{t}=v\\ 0,&\text{ if }m_{t}\in V\setminus\{v\},\end{dcases}$ (32)
@@
    For a sufficiently large $\text{For a sufficiently large $N_{batch}$:}\quad\Pr(m=\xi\mid x,\phi)\approx\frac{|\{x\in\mathcal{B}\mid S_{\phi}(x)=\xi\}|}{N_{batch}}\$ (35)
@@
    =$\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{t=1}^{C}\left(\sum_{v,v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi)\right.$
@@
    (36)${{\approx}}\mathbb{E}_{x,x^{\prime}\in\mathcal{B}}[d_{H}(S_{\phi}(x),S_{\phi}(x^{\prime}))],$ (38)
    In the above step ()($\star$)( ⋆ ) we used that
@@
    Contrastive loss functions such as the triplet loss [52] and NT-Xtent [8] work by defining positive and negative image pairs (usually through image augmentation), and aiming to minimize the distance between embeddings of positive pairs and maximize the distance between negative pairs. Specifically, instead of purposefully sampling negative samples for every $x$, NT-Xtent produces a single pair of samples for each $x$ in a mini-batch $N$ by augmenting each image. Each image in the batch therefore has $12N-$ 1 negative pairs. They use the cosine similarity $sim(u,v)=u^{T}v/\|u\|\|v\|$ and some temperature value $\tau$ to define the following loss between positive pair embeddings $z_{i}$ and $z_{j}$:
@@
    (Jensen $\overset{\begin{subarray}{c}\text{(Jensen}\\ \text{ inequality)}\end{subarray}}{\leq}$ $-1+N-\log N-\frac{1}{N}\sum_{k=1}^{N}d_{H}(m_{i},m_{k}).$
@@ G.4 Straight-Trough Gumbel-Softmax (STGS) trick
As in [18] and subsequent works, the optimization of policy $\pi_{\varphi}$ (minimization over $\varphi$) can in theory be done via policy gradient optimization methods such as REINFORCE [58], however, due to the combinatorial explosion of the number of messages this can be very compute-intensive if calculated directly, or produces high variance estimators. We thus follow the straight-trough Gumbel-softmax trick [19], in which the probabilistic interpretation of $\pi_{\varphi}$ is implicit in the network architecture, and the architecture itself is deterministic and directly differentiable. In this case, we can thus pass to the deterministic description as $m=S_{\phi}(x)$, replacing the probabilistic one $m\sim\pi_{\varphi}(\cdot|m)$, and we obtain the updated form of (47) as
@@
Appendix H Hyperparameter search
Table 3: Comparison of TopSim, Final MSE, and Useful Length for different regularizations and values of $\beta$ and $\lambda$.
Dataset Mode Lambda Beta TopSim Final MSE Useful Length
Shapes3D PDM 1.0 1.25 0.264 $\pm$ 0.010 0.208 $\pm$ 0.020 7.600 $\pm$ 0.780
PDM 1.0 1.5 0.295 $\pm$ 0.010 0.182 $\pm$ 0.020 6.600 $\pm$ 0.610
PDM 5.0 1.25 0.262 $\pm$ 0.020 0.215 $\pm$ 0.030 7.600 $\pm$ 0.780
PDM 5.0 1.5 0.260 $\pm$ 0.010 0.178 $\pm$ 0.010 5.000 $\pm$ 0.280
PDM 10.0 1.25 0.279 $\pm$ 0.010 0.191 $\pm$ 0.020 6.200 $\pm$ 0.660
PDM 10.0 1.5 0.276 $\pm$ 0.010 0.177 $\pm$ 0.010 6.400 $\pm$ 0.610
KoLeo 1.0 1.25 0.222 $\pm$ 0.020 0.272 $\pm$ 0.040 8.200 $\pm$ 0.720
KoLeo 1.0 1.5 0.265 $\pm$ 0.000 0.184 $\pm$ 0.010 7.000 $\pm$ 0.630
KoLeo 5.0 1.25 0.258 $\pm$ 0.010 0.198 $\pm$ 0.000 7.800 $\pm$ 0.660
KoLeo 5.0 1.5 0.247 $\pm$ 0.020 0.222 $\pm$ 0.020 8.000 $\pm$ 0.570
KoLeo 10.0 1.25 0.244 $\pm$ 0.010 0.232 $\pm$ 0.030 7.500 $\pm$ 0.550
KoLeo 10.0 1.5 0.257 $\pm$ 0.010 0.202 $\pm$ 0.010 6.800 $\pm$ 0.500
MPI3D PDM 1.0 1.25 0.154 $\pm$ 0.000 0.020 $\pm$ 0.000 6.200 $\pm$ 1.530
PDM 1.0 1.5 0.160 $\pm$ 0.000 0.020 $\pm$ 0.000 6.600 $\pm$ 1.310
PDM 5.0 1.25 0.152 $\pm$ 0.000 0.020 $\pm$ 0.000 5.000 $\pm$ 1.470
PDM 5.0 1.5 0.158 $\pm$ 0.010 0.020 $\pm$ 0.000 6.800 $\pm$ 1.210
PDM 10.0 1.25 0.161 $\pm$ 0.000 0.020 $\pm$ 0.000 6.200 $\pm$ 1.530
PDM 10.0 1.5 0.162 $\pm$ 0.000 0.020 $\pm$ 0.000 7.800 $\pm$ 1.070
KoLeo 1.0 1.25 0.162 $\pm$ 0.000 0.020 $\pm$ 0.000 4.000 $\pm$ 1.170
KoLeo 1.0 1.5 0.160 $\pm$ 0.000 0.020 $\pm$ 0.000 7.600 $\pm$ 1.250
KoLeo 5.0 1.25 0.155 $\pm$ 0.000 0.020 $\pm$ 0.000 6.600 $\pm$ 1.340
KoLeo 5.0 1.5 0.163 $\pm$ 0.000 0.020 $\pm$ 0.000 5.800 $\pm$ 1.210
KoLeo 10.0 1.25 0.160 $\pm$ 0.000 0.020 $\pm$ 0.000 4.000 $\pm$ 1.170
KoLeo 10.0 1.5 0.159 $\pm$ 0.000 0.020 $\pm$ 0.000 5.800 $\pm$ 1.180
Appendix I Generation examples
I.1 Reconstruction
The following reconstruction examples were generated with the CELEBI model, with PDM regularization, $\beta=1$ and $\lambda=1.5$. From left to right, the first image of each example is the original image to be reconstructed, and the following images are the predicted reconstructions $x_{i}=R_{\omega}(S_{\phi}(x)_{[1:i]})$ with the VAE backbone.
We observe semantic differences between each prediction (floor hue, shape, object color, angle, etc.), as opposed to increasing in image quality or clarity. This fact leads us to believe that the message space effectively captures the generating factors $\mathcal{G}$ and is likely to be compositional in some measure.

### v4

@@ Appendix E Proof that final state reconstruction allows for wider message choice freedom
Our main result here is the following:
@@
1. For $i<C$ and any choice of $M,\epsilon>0$, the restrictions on $i$-token sub-messages of $S(x)$ by requiring $S\in\mathcal{S}_{full}^{M}$ are more restrictive than those imposed by requiring $S\in\mathcal{S}_{final}^{\epsilon}$, i.e. we have $\mathcal{R}_{[i]}(\mathcal{S}_{full}^{M})\subset\mathcal{R}_{[i]}(\mathcal{S}_{final}^{\epsilon})$.
@@
Item 1 is direct, as no restriction on intermediate sub-messages $S(x)_{[i]}=\mathcal{R}_{[i]}(S(x))$ by the requirement $S\in\mathcal{S}_{final}^{\epsilon}$.
@@
The freedom of exploration ensured by item 1 of the proposition is our main motivation: by only constraining the reconstruction of the full message, we allow much more freedom for the encoding strategy of $S(x)$ at earlier sub-messages at each iteration.
Item 2 of Prop. E.1 shows that the two losses in the definition of $\mathcal{S}_{full},\mathcal{S}_{final}$ give comparable guarantees. At the beginning of iterations we will need to use the general bound with $\epsilon/C$, and as $R^{\star}$ will have been trained on more and more messages during the iteration, the hypothesis of $\mathbb{E}_{x}[d_{i}^{S}(x)]$ non-increasing will become true allowing for the sharper bound with the same value of $\epsilon$, as $R^{\star}$ (an LSTM taking as input successive message tokens) will become efficient in using sub-messages of pairs $(m,m^{\prime})=(S(x),S^{\star}(x))$ to gradually distill information allowing to distinguish them for a large set of $S,x$.
@@ Appendix F Regularization to counteract Holistic Encodings
This corresponds to the balance between compositional and parsimonious communication, a general theme in classical IL literature (see e.g. [Kirby2002, KIRBY201587]). It seems useful to specify more precise definitions in our generative setting, in order to help justify our architecture choices.
@@
    In §F.1 we set rigorous definitions of Holistic Encodings within the setting of §2.1 and §$2.$
@@ F.1 Definitions of Holistic and Compositional encodings, and role of the compressibility condition
In the classical IL setup [Kirby2002, KIRBY201587], a holistic system of communication is defined as an encoding $\mathcal{G}\to\mathcal{M}$ that maps $\mathcal{G}\ni\mathbf{G}\mapsto\mathbf{G^{\prime}}\in\mathcal{M}$ injectively but disregards the compositional structure of the "language" to be encoded, i.e. in our case, without respecting the structure of $\mathbf{G}$. In our generation task formalization, a difference is that the relevant encoding map now factors through the generation map as $\mathcal{G}\stackrel{{\scriptstyle\mathsf{GenX}}}{{\longrightarrow}}\mathcal{X}\stackrel{{\scriptstyle S}}{{\longrightarrow}}\mathcal{M}$. Now the generation map $\mathsf{GenX}$ is considered fixed, thus the definition refers to only $S$ and is the following:
@@
Note that, as in [Kirby2002, KIRBY201587], Definition F.1 of holistic encodings refers to "respecting compositional structure", which requires another definition. The definition of compositional encodings crucially depends on how complex we need/want to assume $\mathbf{G}$ to be. In this paper, for the sake of concreteness, we restricted to the simple case of $\mathbf{G}$ being formed as an element of a product space of finite cardinality, i.e., an $n$-ple of independent factors $\mathbf{G}=[G_{1},\dots,G_{n}]$ in which each $G_{i}$ can take finitely many values. We thus use a straightforward toy definition of compositional encodings, requiring that each of the $G_{i}$ is separately encoded as a string and these strings are concatenated together to give the message that encodes $\mathbf{G}$:
@@ F.3 Encodings that optimize efficiency under expressivity over $\mathcal{D}_{train}$ are holistic
    and thus $\log_{|V|}p\leq-1$, and thus the second inequality $(\star\star)$ in (23) is guaranteed to be sharp, showing the desired strict inequality $C_{0}>C_{\star}$ independently of the sizes of factor ranges $N_{i},i=1,\dots,n$.
@@
    then $(\star)$ becomes sharp, again guaranteeing $C_{0}>C_{\star}$ as desired. This could be probable to happen in our setting, and becomes more likely for large $|V|>1/p$ for which (24) fails, given that the $N_{i}$ are supposed to be unknown, so it is likely that for large $|V|>1/p$ for a few of the $N_{i}$ the gaps in (25) are nontrivial. For example, (25) holds true for $|V|>1/p\geq 10$ if, say, at least two of the $G_{i}$ are binary factors, so that $N_{i}=2$.
@@ F.4 Possible choices of encoded message regularization
    Background and standard approach. Introduced by [entropymax] for use in policy-based reinforcement learning, entropy maximization is known to enforce exploration in deterministic policies and avoid early convergence to single output choices for learned policies. Specifically, in a learning setup with input $x$, output $y$ and policy parameters $W$, the authors define the following estimator for fixed output value $\xi$:
@@
    $\ln Pr(y=\xi|W,x)=-\sum_{i\in U}g_{i}(\xi_{i},w^{i},x^{i})=h(y,W,X).$ (30)
    In fact, this method can be extended onto sequence prediction models, and even IL, such as in [ren2020Compositional].
@@
[…]
@@
    $\text{For a sufficiently large $N_{batch}$:}\quad\Pr(m=\xi\mid x,\phi)\approx\frac{|\{x\in\mathcal{B}\mid S_{\phi}(x)=\xi\}|}{N_{batch}}\\$ (34)
@@
    $\overset{\begin{subarray}{c}\text{($-x\ln x\geq x-x^{2}$}\\ \text{for $x\leq 1$)}\end{subarray}}{\geq}\sum_{t=1}^{C}\sum_{v\in V}\left[Pr(m_{t}=v\mid x,\phi)-Pr(m_{t}=v\mid x,\phi)^{2}\right]$
    $\stackrel{{\scriptstyle(\star)}}{{=}}\sum_{t=1}^{C}\left(\sum_{v,v^{\prime}\in V}Pr(m_{t}=v\mid x,\phi)Pr(m_{t}^{\prime}=v^{\prime}\mid x,\phi)\right$.
@@
[…]
    In the above step $(\star)$ we used that
@@
    Contrastive loss functions such as the triplet loss [triplet] and NT-Xtent [chen2020simpleframeworkcontrastivelearning] work by defining positive and negative image pairs (usually through image augmentation), and aiming to minimize the distance between embeddings of positive pairs and maximize the distance between negative pairs. Specifically, instead of purposefully sampling negative samples for every $x$, NT-Xtent produces a single pair of samples for each $x$ in a mini-batch $N$ by augmenting each image. Each image in the batch therefore has 2N12N-1 negative pairs. They use the cosine similarity $sim(u,v)=u^{T}v/\|u\|\|v\|$ and some temperature value $\tau$ to define the following loss between positive pair embeddings $z_{i}$ and $z_{j}$:
@@
    $\overset{\begin{subarray}{c}\text{(Jensen}\\ \text{ inequality)}\end{subarray}}{\leq}$ $-1+N-\log N-\frac{1}{N}\sum_{k=1}^{N}d_{H}(m_{i},m_{k}).$
@@ G.4 Straight-Trough Gumbel-Softmax (STGS) trick
As in [havrylov2017emergence] and subsequent works, the optimization of policy $\pi_{\varphi}$ (minimization over $\varphi$) can in theory be done via policy gradient optimization methods such as REINFORCE [williams1992simple], however, due to the combinatorial explosion of the number of messages this can be very compute-intensive if calculated directly, or produces high variance estimators. We thus follow the straight-trough Gumbel-softmax trick [jang2017categoricalreparameterizationgumbelsoftmax], in which the probabilistic interpretation of $\pi_{\varphi}$ is implicit in the network architecture, and the architecture itself is deterministic and directly differentiable. In this case, we can thus pass to the deterministic description as $m=S_{\phi}(x)$, replacing the probabilistic one $m\sim\pi_{\varphi}(\cdot|m)$, and we obtain the updated form of (46) as
@@
Appendix H Example Generation: Reconstructions
The following reconstruction, Figure 5, examples were generated with the CELEBI model using PDM regularization with $\beta=1$ and $\lambda=1.5$. In each row, the leftmost image corresponds to the original input, while the subsequent images represent the predicted reconstructions $x_{i}=R_{\omega}(S_{\phi}(x)_{[1:i]})$ obtained from progressively longer message prefixes with the VAE backbone.
We observe that the reconstructions do not merely improve in pixel-level fidelity but rather exhibit semantic refinements across successive steps — variations in floor hue, object color, shape, or viewing angle. This progression suggests that the message space encodes the underlying generative factors $\mathcal{G}$ and supports a degree of compositional structure, where messages carry disentangled information.
Qualitative comparison across methods.
To further assess the representational properties of CELEBI, we qualitatively compare reconstructions across the different methods proposed in this work (see Figures 6 and 7). As illustrated in these figures, CELEBI reaches semantically faithful reconstructions significantly earlier in the decoding sequence than the baseline methods.
Moreover, the differences between reconstructions are more semantically pronounced than in baseline models, which often show subtler variations mainly related to texture or brightness. This qualitative evidence supports our hypothesis that CELEBI facilitates a more structured message space, enabling representations that are not only more compact but also more semantically coherent.
@@ Appendix H Example Generation: Reconstructions
(a) Reconstruction using the baseline model.
(b) Reconstruction using partial decoding.
(c) Reconstruction using final state imitation.
(d) Reconstruction using PD+FiSI+PDM.
Figure 6: Reconstruction for the same example with different methods.
(a) Reconstruction using the baseline model.
(b) Reconstruction using partial decoding.
(c) Reconstruction using final state imitation.
(d) Reconstruction using PD+FiSI+PDM.
Figure 7: Reconstruction for the same example with different methods.
Appendix I Additional Language Metrics
Table 3: Comparison of proposed methods using language variation metrics from [conklin2023compositionality].
Synonymy $\downarrow$ Homonymy $\downarrow$ Freedom $\downarrow$ Entanglement $\downarrow$
Interaction Full message reconstruction, no IL (Baseline) $0.563\pm 0.026$ $0.695\pm 0.015$ $0.569\pm 0.026$ $0.831\pm 0.006$
Progressive Decoding (PD, ours) $0.640\pm 0.022$ $0.737\pm 0.013$ $0.646\pm 0.013$ $0.786\pm 0.007$
Imitation Message imitation (baseline) $0.601\pm 0.039$ $0.714\pm 0.027$ $0.613\pm 0.039$ $0.822\pm 0.010$
Full state imitation $0.597\pm 0.017$ $0.734\pm 0.021$ $0.604\pm 0.018$ $0.824\pm 0.007$
Final-State Imitation (FiSI, ours) $0.633\pm 0.038$ $0.751\pm 0.018$ $0.640\pm 0.038$ $0.842\pm 0.007$
Regularization PD+FiSI+KoLeo $\lambda=1.5$ $0.594\pm 0.027$ $0.739\pm 0.013$ $0.598\pm 0.027$ $0.805\pm 0.007$
PD+FiSI+PDM (ours) $\lambda=1.5$ $0.654\pm 0.013$ $0.753\pm 0.008$ $0.658\pm 0.013$ $0.780\pm 0.008$
(a) Shapes3D
Synonymy $\downarrow$ Homonymy $\downarrow$ Freedom $\downarrow$ Entanglement $\downarrow$
Interaction Full message reconstruction, no IL (Baseline) $0.595\pm 0.053$ $0.597\pm 0.061$ $0.597\pm 0.052$ $0.852\pm 0.021$
Progressive Decoding (PD, ours) $0.484\pm 0.044$ $0.486\pm 0.056$ $0.486\pm 0.044$ $0.786\pm 0.025$
Imitation Message imitation (baseline) $0.367\pm 0.034$ $0.393\pm 0.042$ $0.369\pm 0.034$ $0.837\pm 0.015$
Full state imitation $0.362\pm 0.045$ $0.424\pm 0.052$ $0.364\pm 0.044$ $0.832\pm 0.018$
Final-State Imitation (FiSI, ours) $0.324\pm 0.020$ $0.393\pm 0.018$ $0.328\pm 0.020$ $0.830\pm 0.008$
Regularization PD+FiSI+KoLeo $\lambda=1.5$ $0.365\pm 0.011$ $0.338\pm 0.019$ $0.367\pm 0.011$ $0.763\pm 0.005$
PD+FiSI+PDM (ours) $\lambda=1.5$ $0.347\pm 0.012$ $0.333\pm 0.010$ $0.349\pm 0.013$ $0.756\pm 0.005$
(b) MPI3D
In addition to the metrics presented in the main text of this work, we attempted to measure other metrics for qualitative linguistic variation. We were not able to find many such metrics in the literature, and thus restrict to the ones presented in [conklin2023compositionality]. Importantly, since this reference had a slightly different focus, variation in this context will not refer to the evolution of the language in time, but rather a departure from regularity, which masks an underlying compositional structure.
We study the four measures of variation presented in the paper, namely synonymy, homonymy, word order freedom and entanglement. Synonymy measures the presence of one-to-many mappings between atomic meanings and characters in a position, being minimized when each generating factor value is mapped to a single position and character. Homonymy measures the opposite, i.e., the presence of many-to-one mappings, and is minimized when each character in a position is mapped to a unique generating factor value. Word order freedom refers to the strictness of the mapping between generating factors and positions in the message (for example, if the mapping system always encodes the shape generating factor at the first position). It is minimized if all single generating factors are encoded in the same message position. Similarly, entanglement is minimized when all factors are encoded into unique positions in the message.
Important formular metricas, pero no encaja tan bien. Unico paper que conocemos. No clear patterns, but we argue it is because… An important assumption made by the authors for all these metrics is that meaning should be undivisably encoded into single positions of the message. We believe this scope of regularity is too narrow to capture compositionality-respecting mappings in out setting, as defined in F.2, and does not fit well when the space of messages is greatly larger than the amount of possible states. We found no clear trend or similar behavior across both tested datasets, save a small decrease in entanglement when using PD. This may suggest that the increased pressure for efficiency forces individual generating factors to be uniquely distributed in the message, however we believe there is not sufficient evidence to make such a claim considering the previously discussed pitfalls of the metrics.
To our knowledge there are no available official implementations for these metrics. Our versions can be found in the official repository for this work.
Appendix J Permutation Tests
J.1 Emergent language metrics
To assess the statistical significance of each component added to the Iterated Learning (IL) framework proposed in this work, we conducted permutation tests using the SciPy library [2020SciPy-NMeth]. The test statistic was defined as the difference in means, $\bar{A}-\bar{B}$, where $\bar{A}$ corresponds to the arithmetic mean of the evaluated mode and metric across 10 random seeds. The complete results are presented in Table 5(b).
Across both the Shapes3D and MPI3D datasets, we found that the inclusion of the PD module consistently reduced the useful message length, yielding a test statistic of approximately $2.5$ and a $p$-value below $0.05$. None of the other additions introduced in this work produced a statistically significant decrease in useful length relative to their respective baselines, supporting our hypothesis that PD imposes an efficiency pressure on the emergent language. Moreover, we observed a statistically significant increase in TopSim for the Shapes3D dataset, with a test statistic of $0.026$ and $p<0.05$.
Final-State Imitation had a statistically significant gain in TopSim over both baseline methods on MPI3D, with statistic of $\sim 0.02$ and $p<0.02$. PDM outperformed the KoLeo estimator in both reconstruction error and TopSim on Shapes3D, but achieved no statistically significant increase in MPI3D.
Table 4: Permutation tests for Shapes3D and MPI3D on emergent language metrics.
Metric Mode A Mode B Statistic p-Value
TopSim Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) 0.0263 0.0233
TopSim Full state imitation Final-State Imitation (FiSI; ours) -0.0267 0.1014
TopSim Full state imitation Message imitation (Baseline) -0.0018 0.9332
TopSim Final-State Imitation (FiSI; ours) Message imitation (Baseline) 0.0249 0.2605
TopSim PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ -0.0371 0.0004
Useful Length Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) -2.5 0.0031
Useful Length Full state imitation Final-State Imitation (FiSI; ours) 0.0 1.0
Useful Length Full state imitation Message imitation (Baseline) 0.0 1.0
Useful Length Final-State Imitation (FiSI; ours) Message imitation (Baseline) 0.0 1.0
Useful Length PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ 1.3000 0.0996
Final MSE Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) -0.0218 0.0657
Final MSE Full state imitation Final-State Imitation (FiSI; ours) -0.0130 0.4871
Final MSE Full state imitation Message imitation (Baseline) -0.0201 0.5679
Final MSE Final-State Imitation (FiSI; ours) Message imitation (Baseline) -0.0071 0.8384
Final MSE PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ 0.0343 0.0431
(a) Shapes3D permutation test
Metric Mode A Mode B Statistic p-Value
TopSim Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) 0.0037 0.4416
TopSim Full state imitation Final-State Imitation (FiSI; ours) -0.0193 0.0019
TopSim Full state imitation Message imitation (Baseline) 0.0047 0.4395
TopSim Final-State Imitation (FiSI; ours) Message imitation (Baseline) 0.0240 0.0007
TopSim PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ -0.0060 0.4767
Useful Length Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) -2.7000 0.0345
Useful Length Full state imitation Final-State Imitation (FiSI; ours) 0.0231 1.0
Useful Length Full state imitation Message imitation (Baseline) 0.2231 0.3998
Useful Length Final-State Imitation (FiSI; ours) Message imitation (Baseline) 0.2000 0.5820
Useful Length PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ 0.2222 1.0
Final MSE Progressive Decoding (PD; ours) Full message reconstruction; no IL (Baseline) 0.0011 0.5894
Final MSE Full state imitation Final-State Imitation (FiSI; ours) -0.0018 0.1166
Final MSE Full state imitation Message imitation (Baseline) 0.0019 0.2870
Final MSE Final-State Imitation (FiSI; ours) Message imitation (Baseline) 0.0037 0.0201
Final MSE PD+FiSI+KoLeo $\lambda=1.5$ PD+FiSI+PDM $\lambda=1.5$ -0.000012 0.9626
(b) MPI3D permutation test
J.2 Disentanglement permutation tests
In this section we evaluate the statistical significance of the differences between our model and the discrete baseline on the metrics used in table 2. Similar to the previous section, we conducted permutation tests, using the difference in means as a statistic for 10 random seeds. The complete results are presented in Table 5.
We found a significant difference in accuracy on the Shapes3D dataset, with a statistic of 0.114 and $p<0.01$. We found a significant descrease in RMSE on both datasets, with statistics of $0.227$ and $1.151$ for Shapes3D and MPI3D, respectively, and $p\leq 0.001$. We also observed a significant gain in disentanglement on both datasets, with statistics of $0.003$ and $0.023$ for Shapes3D and MPI3D, respectively, and $p<0.05$ for both.
Table 5: Permutation tests for Shapes3D and MPI3D on disentanglement, accuracy and RMSE using messages from discrete models as data for a linnear probe.
Metric Mode A Mode B Diff(A-B) p-Value
Accuracy VAE+EL VAE+CELEBI -0.1139 0.0018
RMSE VAE+EL VAE+CELEBI 0.2266 0.0010
DCI Disentanglement VAE+EL VAE+CELEBI -0.0034 0.0042
(a) Shapes3D permutation test
Metric Mode A Mode B Diff(A-B) p-Value
Accuracy VAE+EL VAE+CELEBI -0.0128 0.0570
RMSE VAE+EL VAE+CELEBI 1.1512 0.0010
DCI Disentanglement VAE+EL VAE+CELEBI -0.0230 0.0002
(b) MPI3D permutation test